find the equation of a line for this points (4,1) and is perpendicular to the line 2y-3x=4
Answer by checkley71(8403)

FIRST WE FIND THE SLOPE OF THE LINE 2Y-3X=4
2Y=3X-4
Y=3X/2-4/2
Y=3X/2-2 THIS LINE HAS A SLOPE OF 3/2 THEREFORE THE PERPENDICULAR LINE HAS A SLOPE OF -2/3.
THUS THE FORMULAR FOR THIS PERPENDICULAR LINE IS FOUND BY REPLACING X & Y WITH THE POINTS (4,1) WITH A SLOPE OF -2/3 THUS
1=-2/3*4+b
1=-8/3+b
1+8/3=b
b=11/3 THUS THIS EQUATION IS
Y=-2/3X+11/3